# Towards a Multi-Embodied Grasping Agent

### v1

@@ Abstract
Multi-embodiment grasping focuses on developing approaches that exhibit generalist behavior across diverse gripper designs. Existing methods often learn the kinematic structure of the robot implicitly and face challenges due to the difficulty of sourcing the required large-scale data. In this work, we present a data-efficient, flow-based, equivariant grasp synthesis architecture that can handle different gripper types with variable degrees of freedom and successfully exploit the underlying kinematic model, deducing all necessary information solely from the gripper and scene geometry. Unlike previous equivariant grasping methods, we translated all modules from the ground up to JAX and provide a model with batching capabilities over scenes, grippers, and grasps, resulting in smoother learning, improved performance and faster inference time. Our dataset encompasses grippers ranging from humanoid hands to parallel yaw grippers and includes 25,000 scenes and 20 million grasps.
@@
I Introduction
Multi-embodiment grasping methods employ a single model to synthesize grasps for diverse gripper architectures with varying degrees of freedom (DoFs). Since these methods are not tied to one physical system, they adapt more effectively to a broad range of end effectors. To succeed in this setting, approaches must be robust to hyperparameter choices and hidden biases, enabling the integration of data from multiple sources [2, 3]. Recent work has combined diffusion models [4, 5] with architectures that leverage data symmetries through equivariance [6, 7, 8, 9] in robotics. These techniques excel at capturing the multimodal nature of grasping data and currently define the state of the art in robotic grasping. In this paper, we extend these ideas to the multi-embodiment setting and introduce a single method capable of synthesizing grasps for grippers with variable DoFs. Our contributions are threefold. We first introduce an equivariant flow-based method that can synthesize grasps across a wide range of gripper types, including humanoid hands such as Allegro and ShadowHand, as well as parallel jaw grippers such as Panda and VX300, encompassing both rotary and prismatic joint types. We provide a complete implementation in JAX [10], which to our knowledge enables, for the first time in this setting, training and inference across multiple full-scene point clouds, grippers, and grasps in a single update step as a single batch. This capability results in smoother learning for both single-GPU and multi-GPU schemes, significantly faster inference, and lower memory consumption compared to prior approaches. In addition, we release the datasets containing grasps with joint information for five gripper types, together with their JAX-integrated kinematic models, spanning more than 25k scenes. We further provide the data-generation scripts that support both single-object and full-scene grasp generation, facilitating dataset expansion. Finally, we demonstrate that our approach achieves performance on part with the state of the art in both single- and multi-gripper grasp generation settings. Code and dataset are available as open source. [^1] Code and dataset will be made publicly available soon
@@ II Related Work
Grasp detection approaches synthesize grasp poses and configurations for objects and are typically categorized into offline methods, which generate grasps once, and online approaches, which refine grasps continuously during execution. Most work focuses on parallel-jaw grippers [6, 11, 12, 13], usually tied to a single gripper architecture and using large-scale grasp datasets such as ACRONYM [14] and GraspNet-1Billion [15].
@@
Dexterous grasping datasets and algorithms target high-DoF gripper architectures. DexGraspNet 2.0 [16] learns a diffusion model to predict pre-grasp poses with joint configurations for the Leap gripper in cluttered scenes, and Fast-Grasp’D [17] generates multi-finger grasps via differentiable simulation. Dexterous grasping agents employ policies dedicated to high-DoF grippers such as the Leap and ShadowHand; notable examples include UniDexGrasp++ [18], Unified Generative Grasping [19], and AffordDexGrasp [20]. Furthermore, DexGraspVLA [21] demonstrates effective dexterous grasping using pretrained vision and language models, and DexGrasp Anything [22] uses a physics-informed diffusion process to generate dexterous grasps. MultiGripperGrasp [23], BoDex [24], and Get-a-Grip [25] contribute scalable training and evaluation resources for diverse hands, while GenDexGrasp [2] demonstrates object-centric features that transfer across grippers.
Cross-embodiment grasp representations define, for each gripper, a contact intent and treat inverse kinematics as a post-processing task. Prominent examples include Geometric Matching [3], which learns contact maps for each gripper type; $D(R,O)$ Grasp [26], which relates gripper and object point clouds, thereby defining contacts; and RobotFingerPrint [27], which enables transfer between hands through a unified gripper representation, as well as CEDex [28], which uses human-informed contacts to condition a variational autoencoder for diverse grippers. Furthermore, methods that learn gripper-specific encodings [29, 30, 1, 31] predict stable grasps by incorporating the gripper geometry. However, these approaches do not directly learn all DoFs of the gripper and either rely on fixed, predefined configurations [29, 1, 31] or reduce the space through eigengrasps [30]. MADRL [32] applies reinforcement learning to grasping across multiple gripper types but requires specific masking to adapt the model to variable DoFs.
@@
Diffusion- and flow-based generative models learn a transport between an initial uninformative distribution and the target distribution approximated from samples. Work such as [33] learns full robot trajectories, while work in grasping typically focuses on target end-effector grasp poses on the $\mathrm{SE}(3)$ manifold [4, 34, 5, 35, 36, 37, 1].
@@
A surge of large-scale manipulation-skill datasets targets the multi-embodiment problem. RoboMind [38], DROID [39], and AgiBot World Colosseo [40] collect large-scale trajectories for arms and humanoids. All of these datasets rely on heavy task standardization and are incompatible with each other. RoboArena [41] addresses this gap with a distributed real-robot benchmark but, at the time of writing, lacks broad acceptance in the community.
@@
Generalist agents leverage large-scale data and foundation models to produce policies that cover diverse tasks [42, 43, 44, 45, 46]. These methods often struggle with an agent–action representation appropriate for multi-embodiment settings. The work of [47] first attempts to learn a world model from images and extract a multi-embodiment policy via an embodiment-agnostic optical flow that is fine-tuned for a concrete embodiment. While these agents demonstrate impressive results, they often treat grasping as a downstream task.
@@
Locomotion research mirrors the multi-embodiment trend. Work such as One Policy to Run Them All [48], Multi-Loco [49], and Towards Embodiment Scaling Laws in Robot Locomotion [50] demonstrates that increasing morphology diversity improves zero-shot transfer to unknown morphologies.
@@
Zero-shot domain transfer methods bridge different data distributions through alignment of source and target domains, as presented in CycleGAN [51]. Notable examples in robotics include Mirage [52] and Shadow [53], which adapt between domains but are limited to a single physical gripper model, introducing additional biases.
@@
Exploiting symmetries in data can improve robustness, data efficiency, and generalization. Work by Cohen and Welling [54] introduced group-equivariant CNNs for images. Subsequent methods extended these ideas to continuous groups [55, 56, 57], including the Euclidean $E(N)$ group, which has proven effective in robotics [6, 37, 5, 7] and imitation learning [8, 34, 9]. Often, equivariant layers are integrated into graph network architectures [58, 59, 5], producing features that capture local and global structure in robotic scenes.
@@
Finally, modern depth sensors and foundation stereo models reduce hardware assumptions for point-cloud pipelines. Stereo-matching methods such as FoundationStereo [60] and the scalable $S^{2}M^{2}$ model [61] provide reliable depth estimation, reinforcing the feasibility of pure point-cloud models.
@@ III-A Flows for Distribution Modeling in Robotics
Continuous Normalizing Flows (CNFs) provide a framework for learning complex probability distributions on smooth manifolds. For robotics applications, we are often interested in the manifold $\mathcal{M}=\mathrm{SE}(3)\times\mathbb{R}^{D}$. A state is given by $T=(R,\mathbf{p},\mathbf{q})$ where, $R\in\mathrm{SO}(3)$ represents an orientation, $\mathbf{p}\in\mathbb{R}^{3}$ a position, and $\mathbf{q}\in\mathbb{R}^{D}$ a general configuration vector.
CNFs learn to transport samples from a simple prior distribution—typically uniform on compact manifolds like $\mathrm{SO}(3)$ and normal on open manifolds like $\mathbb{R}^{3}$ and $\mathbb{R}^{D}$—to a complex target data distribution $p(T)$. This transport is governed by time-dependent vector fields parameterized by a network with learnable weights $\theta$
$\boldsymbol{\omega}_{\theta}$ $:[0,1]\times\mathcal{M}\to\mathbb{R}^{3},$
$\mathbf{v}_{\theta}$ $:[0,1]\times\mathcal{M}\to\mathbb{R}^{3},$
$\dot{\mathbf{q}}_{\theta}$ $:[0,1]\times\mathcal{M}\to\mathbb{R}^{D}.$
These vector fields define velocities, with their outputs residing in the respective tangent spaces $T_{R}\mathrm{SO}(3)$, $T_{\mathbf{p}}\mathbb{R}^{3}$, and $T_{\mathbf{q}}\mathbb{R}^{D}$. The canonical isomorphism for the tangent space of $\mathrm{SO}(3)$ is defined via a skew-symmetric matrix basis, as in [37]. Within the Flow Matching framework, these velocity fields are learned by minimizing the mean-squared error between the predicted flows and ground-truth flows constructed from pairs of samples drawn from the prior and target distributions.
@@ III-B Equivariant Representations
A function $f:X\to Y$ is equivariant if transforming its input is equivalent to transforming its output; formally, $f(D_{X}(r)x)=D_{Y}(r)f(x)$ for any group action $r\in\mathrm{SE}(3)$ with corresponding representations $D_{X}$ and $D_{Y}$. The network layers in this work are designed to be $\mathrm{SO}(3)$-equivariant and translation-invariant. Features are represented as a direct sum of irreducible representations (irreps) of $\mathrm{SO}(3)$, indexed by type $\ell\in\mathbb{N}_{0}$: $\mathbf{f}=\bigoplus_{\ell=0}^{\ell_{\max}}\mathbf{f}^{\ell}$. Type $\ell=0$ features are scalars, $\ell=1$ are vectors, and higher types capture finer angular detail. A rotation $R\in\mathrm{SO}(3)$ acts on these features via the Wigner-D matrices, $D^{\ell}(R)$. The primary mechanism for equivariant interaction is the tensor product denoted as ($\otimes$). For two features $\mathbf{x}^{\ell_{1}}$ and $\mathbf{y}^{\ell_{2}}$, the output of type $\ell$ is computed via Clebsch-Gordan coefficients (Wigner 3j-symbols)
@@
We leverage two variants of the tensor product. The fully connected tensor product (FCTP) introduces a learnable weight for each valid path in Equation 1, enabling it to capture rich rotational correlations between features. When feature interactions are conditioned on a specific direction, we can use a more efficient $\mathrm{SO}(2)$-equivariant tensor product. This is achieved by rotating the local coordinate system to align with the direction vector, which simplifies the CG coefficients and renders the operation sparse [62]. This technique significantly reduces computational and memory costs and is empirically well-suited for encoding directional information. These operations are foundational to modern architectures like EquiformerV2 [63], which we utilize in our network. Geometric and other scalar information is incorporated by modulating the feature vectors. We define a directional modulation operation based on a relative vector $\mathbf{r}_{ji}=\mathbf{r}_{j}-\mathbf{r}_{i}$ and a set of arbitrary scalars $s$
@@ IV Problem Formulation
We consider an arbitrary gripper $g$ with an internal joint configuration $\mathbf{q}_{g}\in\mathbb{R}^{D_{g}}$, where $D_{g}$ denotes its number of degrees of freedom (DoFs), such as $D_{g}=2$ for a parallel-jaw gripper or $D_{g}=22$ for a dexterous hand. The gripper is modeled as a three-dimensional manifold $\mathcal{M}_{g}$ with known geometry. Its forward kinematics are described by the function
$\mathrm{kin}_{g}\colon\mathbb{R}^{D_{g}}\times\{1,\dots,D_{g}\}\to\mathrm{SE}(3),$
which maps a joint configuration $(\mathbf{q}_{g},i)$ to the transform $T_{i}~=~(R_{i},\mathbf{p}_{i})$ of the $i$-th joint relative to the gripper’s base frame. A scene is composed of graspable objects and potential obstacles, observed as a point cloud $\mathbf{o}_{s}$ in a global coordinate frame. The objective is to learn the conditional distribution defined through
$p(T_{g},\mathbf{q}_{g}\mid\mathcal{M}_{g},\mathbf{o}_{s}),$
where the model leverages the known kinematics $\mathrm{kin}_{g}$. Here, $T_{g}\in\mathrm{SE}(3)$ represents the desired pre-grasp pose of the gripper in the scene frame, and $\mathbf{q}_{g}\in\mathbb{R}^{D_{g}}$ specifies the corresponding pre-grasp joint configuration. A grasp, defined by the pair $(T_{g},\mathbf{q}_{g})$, is considered successful if, after positioning the gripper at this pre-grasp state, a subsequent pre-defined closing motion results in a stable configuration that ensures secure manipulation of an object.
(a) VX300s(b) Franka(c) DEX-EE(d) Allegro(e) Shadow
@@ V Method
Our pipeline begins with a geometric scene encoder that converts the raw point-cloud observations into a feature map that represents the geometry. For each process time step, the kinematics encoder encodes the gripper configuration only through the transform information of the forward kinematics defining the joint queries. In the Multiscale Tensor Field the joint query extract grasp relevant scene information and relate the relative pose of the gripper and scene. In the Flow head, pose and joint gradients are decoded for the flow process. Figure 1 provides an overview of our pipeline. Our method is inspired by architectures such as [6, 5, 59] but is rebuilt from the ground up in the JAX [10] framework, with parallelization and memory efficiency as the main decision drivers. The resulting architecture allows batching over all grippers used, multiple scenes, and grasps in both single- and multi-GPU training schemes.

### v2

@@ Abstract
Multi-embodiment grasping aims to develop approaches that exhibit generalist behavior across diverse gripper designs. Existing methods often learn the gripper kinematic structure implicitly and face challenges due to the difficulty of sourcing the required large-scale data. In this work, we present a data-efficient, flow-based, equivariant grasp synthesis architecture that handles different gripper types with variable degrees of freedom and exploits the underlying kinematic model, deducing all necessary information solely from gripper and scene geometry. Unlike previous equivariant grasping methods, we implement all modules in JAX and provide batching capabilities over scenes, grippers, and grasps, resulting in smoother learning, improved performance, and faster inference. Our dataset encompasses grippers ranging from humanoid hands to parallel-jaw designs, including 25,000 scenes and 20 million grasps.
I Introduction
Multi-embodiment grasping methods employ a single model to synthesize grasps for diverse gripper architectures with varying degrees of freedom (DoF). Since these methods are not tied to one physical system, they adapt effectively to a broad range of end-effectors. To succeed, approaches must be robust to hyperparameter choices and hidden biases, enabling the integration of data from multiple sources [1, 2]. Recent work has combined diffusion models [3, 4] with architectures leveraging data symmetries through equivariance [5, 6, 7, 8]. These techniques excel at capturing the multimodal nature of grasping data and currently define the state of the art. We extend these ideas to the multi-embodiment setting, introducing a single method capable of synthesizing grasps for grippers with variable DoF.
Our main contributions are: We introduce an equivariant flow-based method for synthesizing grasps across a wide range of grippers, including humanoid hands (e.g., Allegro, Shadow Hand) and parallel-jaw grippers (e.g., Franka, ViperX), encompassing both rotary and prismatic joints. We provide a complete JAX [9] implementation that enables, to our knowledge, for the first time in this setting, fully batched training and inference across scenes, grippers, and grasps. This yields smoother learning, faster inference, and lower memory consumption than prior approaches. Additionally, we release datasets with joint information for five gripper types and their JAX-integrated kinematic models, spanning over 25,000 scenes, along with data-generation scripts for both single-object and full-scene tasks. Finally, we demonstrate that our approach achieves performance on par with the state of the art in both single- and multi-gripper settings. Code and data are available as open source. [^1] Code and dataset will be made publicly available soon.
Figure 1: Equivariant Gripper Embeddings. An initial gripper configuration (a) is represented by a learned feature embedding $z$. After a physical joint rotation $\Delta R$, the gripper is in a new configuration (b). Our method ensures the features are correspondingly transformed via the Wigner-D matrices, $z^{\prime}=\mathbf{D}(\Delta R)z$, keeping the representation consistent with the physical state.
@@ II Related Work
Grasp detection approaches synthesize grasp poses and configurations, typically categorized into offline methods (generating grasps once) and online approaches (refining continuously during execution). Most research focuses on parallel-jaw grippers [5, 11, 12, 13], usually tied to single architectures and utilizing large-scale datasets like ACRONYM [14] and GraspNet-1Billion [15]. Dexterous grasping research targets high-DoF architectures. DexGraspNet 2.0 [16] learns a diffusion model to predict pre-grasp poses with joint configurations for the Leap gripper in cluttered scenes, while Fast-Grasp’D [17] generates multi-finger grasps via differentiable simulation. Dexterous agents employ policies for high-DoF grippers such as the Leap and Shadow Hand; examples include UniDexGrasp++ [18], Unified Generative Grasping [19], and AffordDexGrasp [20]. Furthermore, DexGraspVLA [21] demonstrates effective grasping using pretrained vision-language models, and DexGrasp Anything [22] employs a physics-informed diffusion process. MultiGripperGrasp [23], BoDex [24], and Get-a-Grip [25] contribute scalable training resources for diverse hands, while GenDexGrasp [1] demonstrates object-centric features transferring across grippers. Cross-embodiment representations define contact intent per gripper, treating inverse kinematics as post-processing. Prominent examples include Geometric Matching [2], learning contact maps per gripper; $D(R,O)$ Grasp [26], relating gripper and object clouds to define contacts; RobotFingerPrint [27], enabling transfer through unified representations; and CEDex [28], utilizing human-informed contacts for conditioning. While methods learning gripper-specific encodings [29, 30, 10, 31] predict stable grasps by incorporating gripper geometry, they do not directly learn all DoFs, relying on fixed configurations [29, 10, 31] or eigengrasp reductions [30]. MADRL [32] applies reinforcement learning across multiple grippers but requires specific masking for variable DoF adaptation.
@@
Diffusion- and flow-based generative models learn a transport between an initial uninformative distribution and a target distribution approximated from samples. While methods such as [33] learn full robot trajectories, work in grasping typically focuses on target end-effector poses on the $\mathrm{SE}(3)$ manifold [3, 34, 4, 35, 36, 37, 10]. Protein generative models, which often serve as a source of inspiration, share a similar objective and also operate on the same manifold [38, 39, 40].
@@
A surge of large-scale manipulation datasets targets the multi-embodiment problem. RoboMind [41], DROID [42], and AgiBot World Colosseo [43] collect trajectories for arms and humanoids. Relying on heavy task standardization, these datasets are mutually incompatible.
@@
Generalist agents leverage large-scale data and foundation models to produce policies covering diverse tasks [44, 45, 46, 47, 48]. These methods often struggle with agent-action representations appropriate for multi-embodiment settings. Wang et al. [49] attempt to learn a world model from images, extracting a policy via embodiment-agnostic optical flow fine-tuned for specific embodiments. While demonstrating impressive results, these agents often treat grasping as a downstream task.
@@
Locomotion research mirrors the multi-embodiment trend. Works such as One Policy to Run Them All [50], Multi-Loco [51], and Towards Embodiment Scaling Laws in Robot Locomotion [52] demonstrate that increasing morphological diversity improves zero-shot transfer to unknown robots.
@@
Zero-shot domain transfer methods bridge distributions by aligning source and target domains [53]. In robotics, Mirage [54] and Shadow [55] adapt between domains but are limited to single gripper models, introducing biases.
@@
Exploiting symmetries improves robustness and efficiency. Cohen and Welling [56] introduced group-equivariant CNNs for images. Subsequent methods extended this to continuous groups [57, 58, 59], including the Euclidean $E(N)$ group, effective in robotics [5, 37, 4, 6] and imitation learning [7, 34, 8]. Equivariant layers are often integrated into graph networks [60, 61, 4] to capture local and global structure.
@@
Finally, modern foundation stereo models reduce hardware constraints for point-cloud pipelines. Methods such as FoundationStereo [62] and the scene reconstruction transformer VGGT [63] provide reliable depth estimation, reinforcing the feasibility of pure point-cloud models.
@@ III-A Flows for Distribution Modeling in Robotics
Continuous Normalizing Flows (CNFs) provide a framework for learning complex probability distributions on smooth manifolds. For robotics applications, we are often interested in the manifold $\mathcal{M}=\mathrm{SE}(3)\times\mathbb{R}^{D}$. A state is given by $T=(R,\mathbf{p},\mathbf{q})$ where $R\in\mathrm{SO}(3)$ represents orientation, $\mathbf{p}\in\mathbb{R}^{3}$ position, and $\mathbf{q}\in\mathbb{R}^{D}$ a general configuration vector.
CNFs learn to transport samples from a simple prior distribution—typically uniform on compact manifolds such as $\mathrm{SO}(3)$ and normal on open manifolds e.g., $\mathbb{R}^{3}$—to a complex target data distribution $p(T)$. This transport is governed by time-dependent vector fields parameterized by a network with learnable weights $\theta$, defined as
$\boldsymbol{\omega}_{\theta},\mathbf{v}_{\theta}:[0,1]\times\mathcal{M}\to\mathbb{R}^{3},\quad\dot{\mathbf{q}}_{\theta}:[0,1]\times\mathcal{M}\to\mathbb{R}^{D},$
where $\boldsymbol{\omega},\mathbf{v}$ and $\dot{\mathbf{q}}$ are the rotation, position and joint vector fields, respectively. These vector fields define velocities, with their outputs residing in the respective tangent spaces $T_{R}\mathrm{SO}(3)$, $T_{\mathbf{p}}\mathbb{R}^{3}$, and $T_{\mathbf{q}}\mathbb{R}^{D}$. The canonical isomorphism for the tangent space of $\mathrm{SO}(3)$ is defined via a skew-symmetric matrix basis [37]. Within the Flow Matching framework, these velocity fields are learned by minimizing the mean-squared error between the predicted flows and ground-truth flows constructed from pairs of samples drawn from the prior and target distributions [38, 37, 40].
@@ III-B Equivariant Representations
A function $f:X\to Y$ is $\mathrm{SE(3)}$-equivariant if transforming its input is equivalent to transforming its output; formally, $f(D_{X}(r)x)=D_{Y}(r)f(x)$ for any group action $r\in\mathrm{SE}(3)$ with corresponding representations $D_{X}$ and $D_{Y}$. The network layers in this work are designed to be $\mathrm{SO}(3)$-equivariant and translation-invariant. Features are represented as a direct sum of irreducible representations (irreps) of $\mathrm{SO}(3)$, indexed by type $\ell\in\mathbb{N}_{0}$: $\mathbf{f}=\bigoplus_{\ell=0}^{\ell_{\max}}\mathbf{f}^{\ell}$. Type $\ell=0$ features are scalars, $\ell=1$ are vectors, and higher types capture finer angular detail. A rotation $R\in\mathrm{SO}(3)$ acts on these features via the Wigner-D matrices, $D^{\ell}(R)$. The primary mechanism for equivariant interaction is the tensor product, denoted by $\otimes$. For two features $\mathbf{x}^{\ell_{1}}$ and $\mathbf{y}^{\ell_{2}}$, the output of type $\ell$ is computed via Clebsch-Gordan coefficients (CG Wigner 3j-symbols) as
@@
We leverage two variants of the tensor product. The fully connected tensor product (FCTP) introduces a learnable weight for each valid path in Equation 1, enabling it to capture rich rotational correlations between features. When feature interactions are conditioned on a specific direction, we can use a more efficient $\mathrm{SO}(2)$-equivariant tensor product. This is achieved by rotating the local coordinate system to align with the direction vector, which simplifies the CG coefficients and renders the operation sparse [64]. This technique significantly reduces computational and memory costs and is empirically well-suited for encoding directional information. These operations are foundational to modern architectures like EquiformerV2 [65]. Geometric and other scalar information is incorporated by modulating the feature vectors. We define a directional modulation operation based on a relative vector $\mathbf{r}_{ji}=\mathbf{r}_{j}-\mathbf{r}_{i}$ and a set of arbitrary scalars $s$, formulated as
@@ IV Problem Formulation
We consider an arbitrary gripper $g$ with an internal joint configuration $\mathbf{q}_{g}\in\mathbb{R}^{D_{g}}$, where $D_{g}$ denotes its DoF, such as $D_{g}=2$ for a parallel-jaw gripper or $D_{g}=22$ for a dexterous hand. The gripper is modeled as a three-dimensional manifold $\mathcal{M}_{g}$ with known geometry. Its forward kinematics are described by the function $\mathrm{kin}_{g}\colon\mathbb{R}^{D_{g}}\times\{1,\dots,D_{g}\}\to\mathrm{SE}(3),$ which maps a joint configuration $(\mathbf{q}_{g},i)$ to the transform $T_{i}~=~(R_{i},\mathbf{p}_{i})$ of the $i$-th joint relative to the gripper’s base frame. A scene is composed of graspable objects and potential obstacles, observed as a point cloud $\mathbf{o}_{s}$ in a global coordinate frame. The objective is to learn the conditional distribution defined through $p(T_{g},\mathbf{q}_{g}\mid\mathcal{M}_{g},\mathbf{o}_{s}),$ where the model leverages the known kinematics $\mathrm{kin}_{g}$. Here, $T_{g}\in\mathrm{SE}(3)$ represents the desired pre-grasp pose of the gripper in the scene frame, and $\mathbf{q}_{g}\in\mathbb{R}^{D_{g}}$ specifies the corresponding pre-grasp joint configuration. A grasp, defined by the pair $(T_{g},\mathbf{q}_{g})$, is successful if, after positioning the gripper at this pre-grasp state, a subsequent pre-defined closing motion results in a stable configuration that ensures secure object manipulation.
(a) ViperX(b) Franka(c) DEX-EE(d) Allegro(e) Shadow
@@ V Method
Our pipeline begins with a geometric scene encoder that converts raw point-cloud observations into a feature map representing the geometry. For each time step, the kinematics encoder encodes the gripper configuration only through transform information of the forward kinematics defining joint queries. In the Multiscale Tensor Field, joint queries extract grasp-relevant scene information and relate the relative pose of the gripper and scene. In the Flow head, pose and joint gradients are decoded for the flow process. Figure 2 provides an overview of our pipeline. Our method is inspired by architectures such as [5, 4, 61] but is rebuilt from the ground up in the JAX [9] framework, with parallelization and memory efficiency as the main decision drivers. The resulting architecture allows batching over all grippers, multiple scenes, and grasps in both single- and multi-GPU training schemes.